OpenStudy (anonymous):

What is the value of the system determinant for the following system of equations? 6x + y = 4 5x + y - 7 = 0

OpenStudy (anonymous):

-1 1 31 are the answer choices

OpenStudy (anonymous):

6x +y =4 5x+y-7 =0 Move -7 to the right side of "=" 6x +y =4 5x +y =7 Now put it to your calculator. mine is casio fx-115ES If you use the same calculator as mine, u r lucky. Chose mode > 5 > 1 into EQN mode put in 6, 1, 4, 5, 1, 7. then you will have an answer as x= -3, y=22 Test it by plug it into your question and see if they are right

OpenStudy (anonymous):

What is the value of the system determinant for the following system of equations? 6x + y = 4 5x + y - 7 = 0 or 5x+y=7. The determinant for this system of equations is 6*1-5*1=1. I hope the above steps were helpful.